OpenStudy (keana):

9r-2(r=3)=5(3r-4)-26 A: r= ? B: every real number is a solution C: the equation has no solution

Nnesha (nnesha):

two `=` equal signs ?

OpenStudy (keana):

9r-2(r+3)=5(3r-4)-26 my bad

Nnesha (nnesha):

\[\huge\rm 9r-2(r+3)=5(3r-4)-26\] is this ur quesiton ?

OpenStudy (keana):

yeah

Nnesha (nnesha):

oh alright thats fine 1st) distribute parentheses by outside term

Nnesha (nnesha):

\[\huge\rm 9r\color{blue}{-2}(r+3)=\color{ReD}{5}(3r-4)-26\] distribute (3r-4) by 5 and distribute (r+3) by -2

Nnesha (nnesha):

familiar with the distributive property ?

OpenStudy (keana):

9r-2r-6=15r-20--26?

Nnesha (nnesha):

nice combine like terms!

Nnesha (nnesha):

it's -26

OpenStudy (keana):

7r-6=15r+(-46)?

OpenStudy (keana):

-6=8r+(-46)?

Nnesha (nnesha):

right!

OpenStudy (keana):

now what?

Nnesha (nnesha):

now move the numbers to the other side of the equal sign

Nnesha (nnesha):

-46 to the left side then solve for r

OpenStudy (keana):

-40=8r r=-5

Nnesha (nnesha):

you should take the variables term on one side and number to the opposite side

Nnesha (nnesha):

nvm. that's right!

OpenStudy (keana):

so is -5 the answer

Nnesha (nnesha):

wait a sec i have to check it1 more time

OpenStudy (keana):

i tried -5 and it said it was wrong so that cannot be right lol

Nnesha (nnesha):

would you add or subtract 46 both sides ?

Nnesha (nnesha):

it's already negative on the right side how would you cancel out -46 from right side ?

OpenStudy (keana):

i added 46 to each side

Nnesha (nnesha):

yes right so -6+46= ???

Nnesha (nnesha):

46 id greater than -6 so sign would be what as a final answer ?

OpenStudy (keana):

i got -40

Nnesha (nnesha):

no

OpenStudy (keana):

-6+46 is -40

Nnesha (nnesha):

no

Nnesha (nnesha):

if bigger number is negative `then` answer would be negative -3+2=-1 3-2=1

Nnesha (nnesha):

do you see the difference ?

OpenStudy (keana):

ok

Nnesha (nnesha):

so 46 is greater than 6 right ?

OpenStudy (keana):

yeah

Nnesha (nnesha):

yes so if 46 is positive answer would be positive when you subtract if 46 is negative then would be negative -46+6=-40 `but` 46-6=40

OpenStudy (keana):

ok yeah i made a typo on my calculator

OpenStudy (keana):

so 5?

Nnesha (nnesha):

yes right

Nnesha (nnesha):

you can check ur answer substitute x for 5 if you get equal sides then yes 5 is right !

OpenStudy (keana):

ok it was right, awesome

Nnesha (nnesha):

i mean for r**

OpenStudy (keana):

thanks a ton

Nnesha (nnesha):

yep:=) if you replaced r for -5 you wouldn't get the equal sides

Nnesha (nnesha):

my pleasure!! :=)